1) Find the derivative of 2x - e8r)sin(21)
Answer #1

\text{ let }y =(2x-e^{8x})^{sin(2x)}

\left (\text{ We will calculate the derivative using logarithmic differentiation. } \right )

\text{ Taking log on both sides }

ln(y) =ln \left ( (2x-e^{8x})^{sin(2x)} \right )

=> ln(y) = sin(2x)\:\:\:\: ln \left (2x-e^{8x} \right )

\text{ Differentiating with respect to }x

=> \frac{\mathrm{d} }{\mathrm{d} x}ln(y) = \frac{\mathrm{d} }{\mathrm{d} x}\left (sin(2x)\:\:\:\: ln \left (2x-e^{8x} \right ) \right )

=>\frac{1}{y} \frac{\mathrm{d}y }{\mathrm{d} x}=\left ( \frac{\mathrm{d} }{\mathrm{d} x}(sin(2x) \right ) \:\:\:\: ln \left (2x-e^{8x} \right )+sin(2x) \left ( \frac{\mathrm{d} }{\mathrm{d} x} \left ( ln \left (2x-e^{8x} \right )\right )\right )

=>\frac{1}{y} \frac{\mathrm{d}y }{\mathrm{d} x}=\left ( 2cos(2x) \right ) \:\:\:\: ln \left (2x-e^{8x} \right )+sin(2x) \left (\left ( \frac{1}{2x-e^{8x}} \right )\left ( \frac{\mathrm{d} }{\mathrm{d} x} \left (2x-e^{8x}\right ) \right )\right )

=>\frac{1}{y} \frac{\mathrm{d}y }{\mathrm{d} x}= 2cos(2x) \: ln \left (2x-e^{8x} \right )+sin(2x) \left (\left ( \frac{1}{2x-e^{8x}} \right ) \left (2(1)-8e^{8x}\right ) \right )

=>\frac{1}{y} \frac{\mathrm{d}y }{\mathrm{d} x}= 2cos(2x) \: ln \left (2x-e^{8x} \right )+ \frac{sin(2x) (2-8e^{8x})}{2x-e^{8x}}

=> \frac{\mathrm{d}y }{\mathrm{d} x}=y\left ( 2cos(2x) \: ln \left (2x-e^{8x} \right )+ \frac{sin(2x) (2-8e^{8x})}{2x-e^{8x}} \right )

=> \frac{\mathrm{d}y }{\mathrm{d} x}=(2x-e^{8x})^{sin(2x)}\left ( 2cos(2x) \: ln \left (2x-e^{8x} \right )+ \frac{sin(2x) (2-8e^{8x})}{2x-e^{8x}} \right )

Hence.

\text{ Derivative of }(2x-e^{8x})^{sin(2x)} \text{ is }

=(2x-e^{8x})^{sin(2x)}\left ( 2cos(2x) \: ln \left (2x-e^{8x} \right )+ \frac{sin(2x) (2-8e^{8x})}{2x-e^{8x}} \right )